SOS! I really need help here. thank you very much.
convert to logarithmic equations. For example, the logarithmic form of "23 = 8" is "log2 8 = 3".

16 3/2 = 64

Answer by josmiceli(6783)

16^(3/2) = 64
(16^(1/2))^3 = 64
4^3 = 64
(2^2)^3 = 64
2^6 = 64
log(base 2)64 = 6